OpenStudy (anonymous):

Find the product. y^4 · y^3

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow a^m * a^n = a^{m + n} \Longleftrightarrow y^4 * y^3 = ? }\)

OpenStudy (anonymous):

y^7

Parth (parthkohli):

Correct =P

OpenStudy (unklerhaukus):

\[y^4\cdot y^3\]\[=yyyy\cdot yyy=yyyyyyy=y^7\]